A drag-racing car goes from 0 to 290 mph in 5.2 seconds. What is it’s average acceleration in g’s?
1
Expert's answer
2020-01-31T09:53:16-0500

"a=(v1-v0)\/t=(290-0)\/5.2=55.77 meter\/sec^2=5.7g"